# Gaze on the Prize: Shaping Visual Attention with Return-Guided Contrastive Learning

### v1

@@ Abstract
Keywords Reinforcement Learning $\cdot$ Contrastive Learning $\cdot$ Robotic Manipulation $\cdot$ Gaze
footnotetext: Project page: https://andrewcwlee.github.io/gaze-on-the-prize
@@ 1 Introduction
    Learnable Foveal Attention for RL: We adapt a parametric attention model from human gaze research to visual RL. This design provides an inductive bias that is well-suited for manipulation tasks and provides explainable insights into the agent’s actions.
@@
    Return-guided Attention Learning: We devise a contrastive learning method that shapes attention patterns by comparing how an agent’s gaze influences the task outcome. Using a triplet loss, the attention mechanism learns to focus on image regions that distinguish success from failure.
@@
    Plug-in Compatibility: Our approach serves as a plug-in enhancement that is agnostic to the underlying RL framework. It adds a gaze module to improve visual RL algorithms while preserving their logic and structure.
@@ 2.1 Gaze in Computer Vision and Robotics
In robotics, gaze has been studied in both imitation learning (IL) and reinforcement learning (RL). Early IL approaches used Gaussian mixture models to estimate human gaze, foveating a robot’s vision by cropping its input around the predicted gaze kim2020using . A more recent work leverages foveated vision transformers, allocating high-resolution patches near the gaze and coarser patches toward the periphery chuang2025look . In RL, VisaRL liang2024visarl pretrains a vision encoder using human-annotated saliency maps, improving downstream performance for robot control. Eye-Robot kerr2025eye trains an RL agent to control a mechanical eyeball, with rewards guided by the behavior cloning loss of a co-trained IL policy. In contrast to these approaches, we explore how gaze and attention can emerge naturally within standard visual RL frameworks using contrastive learning, without relying on explicit gaze supervision or specialized hardware.
@@ 2.2 Contrastive Learning in RL
Contrastive learning has emerged as a powerful approach for representation learning in RL. CURL improves sample efficiency in RL by applying an instance-level contrastive loss to augmented image observations alongside the standard training objective laskin2020curl . M-CURL extends CURL by adding a masked contrastive objective over video frame sequences, using a transformer to reconstruct features that match the ground truth zhu2022masked . eysenbach2022contrastive reinterprets goal-conditioned RL as a contrastive learning problem, aligning state-action embeddings to approximate a goal-conditioned value function. TACO introduces a temporal contrastive objective to align state-action features with future state representations zheng2023taco . Most closely related, liu2021return uses reward returns to define a contrastive loss, aligning state-action pairs with similar returns. Our work complements and extends this idea by investigating how return-guided contrastive learning can instead shape visual attention, improving both performance and explainability.
@@ 3.1 Overview
The goal of Gaze on the Prize is to enhance visual RL by constraining visual representation learning to a specific region(s) within an image. This is achieved through an attention mechanism trained with return-guided contrastive learning. Our framework is designed as a versatile plugin, the main additions being a simple gaze module and an auxiliary contrastive loss that enhances the base RL algorithm. It is fully compatible with any vision-based RL architecture including those with standard CNN-based vision encoders.
@@ 3.3.1 Contrastive Attention Loss
Given an anchor observation $o_{a}$, we identify its $k$-nearest neighbors in the feature space and partition them into two groups based on their returns. We partition them into high-return (positive) and low-return (negative) groups, using the 50th percentile ($\tilde{R}$) as the dividing threshold with an adaptive delta ($\Delta$) based on return spread. States with returns above $\tilde{R}+\Delta$ become positive samples $\mathcal{P}(o_{a})$, while those below $\tilde{R}-\Delta$ become negative samples $\mathcal{N}(o_{a})$.
@@ 3.3.2 Contrastive Buffer and Return-guided Triplet Mining
The contrastive loss (Equation 2) requires triplets of similar features with different returns. We first describe our buffer for storing vision features, then detail how we mine the triplets from this buffer.
During training, we maintain a buffer (separate from the on-policy batch or off-policy replay buffer) storing three elements from each observation: (1) detached feature maps from the vision backbone’s final layer, (2) flattened feature embeddings for efficient similarity search, and (3) associated episode returns. This circular buffer stores diverse historical visual features rather than just recent features. The vision features are stored detached from the computation graph for gradient isolation, preventing the contrastive loss from participating in vision backbone training. During training, stored features pass through the current attention head, creating gradient paths that update attention while preserving the learned vision representations. This design also reduces computational cost and memory requirements compared to backpropagating through the full network.
We implement triplet mining using FAISS douze2024faiss ; johnson2019billion for efficient $k$-nearest neighbor search. FAISS enables sub-linear search complexity even with large search space, making our approach practical for long training runs. During each mining iteration, we sample anchor features from the buffer and retrieve their $k$-nearest neighbors based on cosine similarity of L2-normalized features. These neighbors form candidate pools for triplet construction. We partition them into positive (high-return) and negative (low-return) groups and randomly sample one from each group to pair with the anchor, forming triplets for contrastive learning.
@@ 4 Experiments
RQ2: Does return-guided contrastive learning enhance attention? We evaluate foveal attention with and without contrastive learning to isolate the contribution of our contrastive learning.
RQ3: Does contrastive learning improve robustness to visual clutter? We train RL agents in cluttered environments to test whether our attention can learn to filter irrelevant visual information despite distractions.
@@ 4.1 Experimental Setup
4.2 Implementation details
We build on widely used PPO schulman2017proximal and SAC haarnoja2018soft implementations provided in ManiSkill3 and CleanRL huang2022cleanrl . For baselines, we either use the recommended hyperparameters from the benchmark or lightly tuned variants to ensure stable training. We do not perform heavy task-specific tuning, since our goal is to test whether the foveal attention and return-guided contrastive learning provide consistent improvements to the base RL algorithm without altering the base RL training behavior. This ensures that our comparisons reflect the effect of our method rather than differences in baseline optimization. For our foveal attention and return-guided contrastive components, we adopt a single set of default hyperparameters across all tasks: the number of anchors is 1024, the top kk=16 nearest neighbors are used for sampling positives and negatives, the contrastive buffer holds 100,000 samples, the contrastive loss weight of $\lambda_{\text{attn}}$=0.1, and the contrastive objective is applied every iteration. We emphasize that these values were chosen as reasonable defaults and may not be individually optimal for each task, though we include ablation on key hyperparameters to examine their influence on performance. A full hyperparameter table for both RL algorithms and our contrastive components is provided in Appendix B.
@@ 5 Results
Figure 2: PPO results on ManiSkill3 environments. Top: Comparison of attention architectures without contrastive learning (RQ1), showing baseline CNN, patch attention, and foveal attention. Center: Effect of return-guided contrastive learning on foveal attention performance (RQ2), comparing foveal attention with and without contrastive supervision. Bottom: Comparison of foveal attention with and without contrastive learning when trained in the presence of visual clutter (RQ3), demonstrating the robustness benefits of return-guided supervision.
@@ 5.1 How do different attention mechanisms affect visual RL performance? (RQ1)
We evaluate PPO agents on six ManiSkill3 tasks, comparing three architectural variants: (1) a baseline without attention, (2) patch attention, and (3) our foveal attention, all without contrastive learning. This comparison is to isolate the effect of attention architecture independent of the contrastive learning objective. As shown in the top row of Figure 2, both attention mechanisms increase sample efficiency (measured by steps to 50% success rate) in five out of six tasks compared to the baseline without attention. Additionally, we observe that our foveal attention consistently reaches the 50% success rate faster than patch attention across all tasks. These findings suggest that while spatial attention generally adds value when training visual RL, the architectural choice matters. We observe that structured focus provides a better inductive bias for object-centric manipulation, which usually requires focused localization throughout the task. Interestingly, we noticed that on PushT task, patch attention underperforms the baseline, suggesting that unconstrained weights can deteriorate training. Without structural constraints, attention may focus on misleading features or shift too rapidly during training which leads to unstable training. On the other hand, our foveal attention appears to provide essential regularization, preventing these failure modes while maintaining flexibility to focus on task-relevant regions. See Figure 3 for the visualization of the two attention variants on an example task.
@@
5.2 Does return-guided contrastive learning enhance attention? (RQ2)
@@
5.3 Does contrastive learning improve robustness to visual clutter? (RQ3)
To further investigate whether contrastive learning helps the attention mechanism to filter irrelevant visual information, we train and evaluate on cluttered variants of a subset of the tasks. As shown in the bottom row of Figure 2, the performance gap is more apparent. For example, while foveal attention without contrastive learning is unable to solve the PushTClutter task, even underperforming the baseline, contrastive learning provides the necessary guidance to find critical cues from the cluttered environment. Also for PokeCubeClutter, the foveal attention with contrastive learning reaches a robust performance of  90% success rate, while without contrastive learning, the success rate plateaus below 80%.
@@ 5.4 Is our approach applicable across different RL algorithms? (RQ4)
Figure 4: SAC results on ManiSkill3 environments (RQ4). Foveal attention with contrastive learning consistently improves off-policy learning across five manipulation tasks, demonstrating the generalizability of our approach beyond on-policy methods.
@@ 5.5 Ablations
We conduct ablations on the PushT task under PPO, as this setting shows the largest gap between the foveal attention and our full method with contrastive learning. This makes it the most informative testbed for isolating the effects of each design choice of our method. We focus our ablation on three core hyperparameters that characterize the general contrastive learning pipeline: (1) buffer size determines the data regime for contrastive learning, (2) loss weight $\lambda_{\text{attn}}$ controls the balance between auxiliary and the main RL objective, and (3) update frequency affects computational cost. These represent fundamental design decisions that apply across any vision-based RL task, rather than task-specific tuning parameters. Figure 5 illustrate how these hyperparameters affect performance, showing the method’s robustness to these critical implementation choices.
Figure 5: Ablation studies on PushT examining key hyperparameters and contrastive learning contribution. Left: Buffer size comparison including no-contrastive baseline. Center: Contrastive loss weight $\lambda_{\text{attn}}$. Right: Update frequency. Results demonstrate robustness to hyperparameter choices within reasonable ranges.
@@
Buffer size: Contrastive buffer size is a notable design choice of our method, as PPO typically uses only the most recent rollout data unlike off-policy methods. Prior work has shown that augmenting on-policy algorithms with replay buffers can improve performance wang2016sample ; novati2019remember , motivating our use of a buffer for contrastive learning. Interestingly, using only on-policy batch ($\approx$20k) performs comparably to the larger 100k buffer, suggesting that our method adapts effectively to different data availability constraints without requiring extensive historical samples in the buffer. However, we noticed that excessively large buffers (200k, 10x on-policy batch) can slow early learning, likely due to mining triplets from stale representations that no longer reflect the current features, highlighting the importance of maintaining a reasonable buffer size that provides sufficient data for contrastive learning without introducing staleness issues.
Contrastive Loss Weight: Balancing auxiliary objectives with the main RL objective typically requires careful tuning to prevent training instability jaderberg2016reinforcement . Our method exhibits minimal sensitivity to the contrastive loss weight ($\lambda_{\text{attn}}$) with all values from 0.05 to 0.25 achieving similar convergence trajectories and final performance. We observe that higher weights (0.25 vs 0.1) produce steeper learning curves, which suggest stronger contrastive supervision accelerates convergence. However, excessive weights can be detrimental, as can be seen with $\lambda_{\text{attn}}$=0.5, suggesting that beyond a certain threshold, the contrastive objective interferes with policy learning. The trade-offs within the acceptable range are minimal compared to the gain over the baseline without any contrastive objective, which indicates that the return-guided signal is inherently beneficial regardless of precise weighting within a reasonable range. This finding makes the method practical for deployment without excessive tuning.
@@
Table 1: Throughput reduction and performance trade-offs. Steps Per Second (SPS) is measured during training on PushT task using a single NVIDIA RTX 3090 GPU, representing wall-clock throughput. Sample efficiency is measured as the ratio of steps required to reach 50% success rate (SR) compared to baseline. The values are averaged across three seeds.
@@
Contrastive Frequency Since triplet mining and contrastive learning add computational overhead, we analyze different update frequencies jaderberg2016reinforcement to determine whether less frequent updates maintain performance benefits. The contrastive frequency plot shows a non-monotonic relationship where the contrastive update of every 2 iterations (yellow) provides the best trade-off, while update every 8 iterations (red) underperforms the baseline despite the added overhead. This suggests that the attention may deviate from task-relevant patterns if they are not updated frequently enough, emphasizing the necessity of providing consistent corrective signals during training.
The throughput reduction and performance trade-offs are organized in Table 1. The results show that our approach not only improves sample efficiency but also achieves faster wall-time convergence despite the added computational overhead. This, along with other results presented in this work demonstrates that the computational investment in learning visual attention is well-justified, leading to practical speedups for visual RL training.

### v2

@@ 1 Introduction
    Learnable Foveal Attention for Visual RL: We adapt a parametric attention model from human gaze research to visual RL. This design provides an inductive bias that is well-suited for manipulation tasks and provides explainable insights into the agent’s actions.
@@
    Return-guided Attention Learning: We devise a contrastive learning method that shapes attention patterns by comparing how an agent’s gaze influences the task outcome. Using triplet loss, the attention mechanism learns to focus on image regions that distinguish success from failure.
@@
    Plug-in Compatibility: Our approach serves as a plug-in enhancement that is agnostic to the underlying RL framework. It adds a dedicated gaze module to improve visual RL algorithms while preserving their logic and structure.
@@ 2.1 Gaze in Computer Vision and Robotics
In robotics, gaze has been studied in both imitation learning (IL) and reinforcement learning (RL). Early IL approaches used Gaussian mixture models to estimate human gaze, foveating a robot’s vision by cropping its input around the predicted gaze kim2020using. A more recent work leverages foveated vision transformers, allocating high-resolution patches near the gaze and coarser patches toward the periphery chuang2025look. In RL, VisaRL liang2024visarl pretrains a vision encoder using human-annotated saliency maps, improving downstream performance for robot control. Eye-Robot kerr2025eye trains an RL agent to control a mechanical eyeball, with rewards guided by the behavior cloning loss of a co-trained IL policy. In contrast to these approaches, we explore how gaze and attention can emerge naturally within training of standard visual RL frameworks using contrastive learning, without relying on explicit gaze supervision or specialized hardware.
@@ 2.2 Contrastive Learning in RL
Contrastive learning has emerged as a powerful approach for representation learning in RL. CURL improves sample efficiency in RL by applying an instance-level contrastive loss to augmented image observations alongside the standard training objective laskin2020curl. M-CURL extends CURL by adding a masked contrastive objective over video frame sequences, using a transformer to reconstruct features that match the ground truth zhu2022masked. eysenbach2022contrastive reinterprets goal-conditioned RL as a contrastive learning problem, aligning state-action embeddings to approximate a goal-conditioned value function. TACO introduces a temporal contrastive objective to align state-action features with future state representations zheng2023taco. Also related is work by, liu2021return which uses reward returns to define a contrastive loss, aligning state-action pairs with similar returns. Our work complements and extends this idea by investigating how return-guided contrastive learning can instead shape visual attention, improving both performance and explainability.
While our objective and mechanism differ from CURL which aims to learn robust visual representations, both operate at the visual representation level and influence how information is extracted from pixel observations. For this reason, we also include a direct empirical comparison to CURL to position our attention learning approach in relation to the broader family of contrastive representation learning methods.
@@ 3.1 Overview
The goal of Gaze on the Prize is to improve visual RL in manipulation by learning a constrained visual focus to a specific region(s) within a view, similar to how humans focus on task relevant regions while performing tasks. This is achieved through an attention mechanism trained with return-guided contrastive learning. Our framework is designed as a versatile plugin, the main additions being a simple gaze module and an auxiliary contrastive loss that enhances the base RL algorithm. It is fully compatible with any vision-based RL architecture including those with standard CNN-based vision encoders.
@@ 3.3.1 Contrastive Attention Loss
Given an anchor observation $o_{a}$ with return $R_{a}$, we identify its $k$-nearest neighbors in the feature space and partition them into two groups based on their returns. Neighbors with returns greater than the anchor’s return ($R>R_{a}$) form the positive pool $\mathcal{P}(o_{a})$, while those with lower returns ($R<R_{a}$) form the negative pool $\mathcal{N}(o_{a})$. From each pool, we sample from a predefined top percentile for positives and a predefined bottom percentile for negatives, ensuring that sampled negatives always have lower returns than the anchor and sampled positives always have higher returns. This guarantees a clean ordering in return space and yields stable and well-separated triplets.
@@ 3.3.2 Contrastive Buffer and Return-guided Triplet Mining
The contrastive loss (Equation 2) requires triplets of similar features with different returns. We first describe our buffer for storing vision features, then detail how we mine the triplets from this buffer. During training, we maintain a buffer (separate from the on-policy batch or off-policy replay buffer) storing three elements from each observation: (1) detached feature maps from the vision backbone’s final layer, (2) flattened feature embeddings for efficient similarity search, and (3) associated episode returns. These features are extracted directly from the same PPO batches or SAC replay-buffer transitions used for RL training and no additional data collection is required. This circular buffer stores diverse historical visual features rather than just recent features. The vision features are stored detached from the computation graph for gradient isolation, preventing the contrastive loss from participating in vision backbone training. During training, stored features pass through the current attention head, creating gradient paths that update attention while preserving the learned vision representations. This design also reduces computational cost and memory requirements compared to backpropagating through the full network.
We implement triplet mining using FAISS douze2024faiss; johnson2019billion for efficient $k$-nearest neighbor search. FAISS enables sub-linear search complexity even with large search space, making our approach practical for long training runs. A persistent FAISS index is incrementally updated as new samples enter the buffer. During each mining iteration, we sample anchor features from the buffer and retrieve their $k$-nearest neighbors based on cosine similarity of L2-normalized features. These neighbors form candidate pools for triplet construction. We partition them into positive (high-return) and negative (low-return) groups and randomly sample one from each group to pair with the anchor, forming triplets for contrastive learning.
Finally, because our triplet construction relies on distinguishing higher- and lower-return neighbors, the contrastive objective assumes that returns exhibit some degree of variation during training. This requirement is not specific to our formulation but is inherent to return-based contrastive learning in general. For example, RCRL liu2021return constructs positive and negative samples by segmenting trajectories according to their returns, an approach that implicitly relies on return variation to provide informative supervision. However, the return-segmentation used in RCRL assumes that transitions with identical or thresholded returns are behaviorally similar, which is an approximation that becomes unreliable in contact-rich manipulation, where visually and geometrically distinct states often share identical rewards. In contrast, modern manipulation benchmarks such as ManiSkill3 taomaniskill3 and Meta-World mclean2025metaworld naturally provide shaped or semi-dense rewards arising from multi-stage task structure and partial-progress signals, which produce sufficient return diversity for stable triplet construction without requiring additional segmentation and approximation.
@@ 4.1 Research Questions
RQ2: Does return-guided contrastive learning enhance foveal attention? We evaluate foveal attention with and without contrastive learning to isolate the contribution of our contrastive learning.
RQ3: Does return-guided contrastive learning help the agent focus on task-relevant regions in the presence of distractor objects? We train RL agents in cluttered environments to test whether our attention can learn to filter irrelevant visual information despite distractions.
@@
RQ5: How does our approach compare against existing contrastive representation-learning baselines? We compare our method against CURL, a contrastive learning visual RL method, to evaluate how a standard augmentation-based contrastive objective transfers to visually complex manipulation settings and how it differs from our return-guided attention learning framework.
@@ 4.2 Experimental Setup
We build on widely used PPO schulman2017proximal and SAC haarnoja2018soft implementations provided in ManiSkill3 and CleanRL huang2022cleanrl. For baselines, we either use the recommended hyperparameters from the benchmark or lightly tuned variants to ensure stable training. We do not perform heavy task-specific tuning, since our goal is to test whether the foveal attention and return-guided contrastive learning provide consistent improvements to the base RL algorithm without altering the base RL training behavior. This ensures that our comparisons reflect the effect of our method rather than differences in baseline optimization. For our foveal attention and return-guided contrastive components, we adopt a single set of default hyperparameters across all tasks: the number of anchors is 1024, the top kk=16 nearest neighbors are used for sampling positives and negatives, the contrastive buffer holds 100,000 samples, the contrastive loss weight of $\lambda_{\text{attn}}$=0.1, and the contrastive objective is applied every iteration. We emphasize that these values were chosen as reasonable defaults and may not be individually optimal for each task, though we include ablation on key hyperparameters to examine their influence on performance. A full hyperparameter table for both RL algorithms and our contrastive components is provided in Appendix D.
@@ 5 Results
Figure 2: PPO results on ManiSkill3 environments. Top: Comparison of attention architectures without contrastive learning (RQ1), showing baseline CNN, patch attention, and foveal attention. Center: Effect of return-guided contrastive learning on foveal attention performance (RQ2), comparing foveal attention with and without contrastive supervision. Bottom: Comparison of foveal attention with and without contrastive learning when trained in the presence of visual distractor objects (RQ3), demonstrating the robustness benefits of return-guided supervision.
@@ 5.1 How do different attention mechanisms affect visual RL performance? (RQ1)
We evaluate PPO agents on six ManiSkill3 tasks, comparing three architectural variants: (1) a baseline without attention, (2) patch attention, and (3) our foveal attention, all without contrastive learning. This comparison is to isolate the effect of attention architecture independent of the contrastive learning objective. For the foveal attention without contrastive learning, we retain the same parametric Gaussian attention head but remove the return-guided contrastive supervision. Thus, the attention parameters receive gradients only from the RL objective. This ensures that any performance difference is attributable purely to the architectural structure of the attention mechanism rather than the additional contrastive signal. As shown in the top row of Figure 2, both attention mechanisms increase sample efficiency (measured by steps to 50% success rate) in five out of six tasks compared to the baseline without attention. Additionally, we observe that our foveal attention consistently reaches the 50% success rate faster than patch attention across all tasks. These findings suggest that while spatial attention generally adds value when training visual RL, the architectural choice matters. We observe that structured focus provides a better inductive bias for object-centric manipulation, which usually requires focused localization throughout the task. Interestingly, we noticed that on PushT task, patch attention underperforms the baseline, suggesting that unconstrained weights can deteriorate training. Without structural constraints, attention may focus on misleading features or shift too rapidly during training which leads to unstable training. On the other hand, our foveal attention appears to provide essential regularization, preventing these failure modes while maintaining flexibility to focus on task-relevant regions. See Figure 3 for the visualization of the two attention variants on an example task.
@@
5.2 Does return-guided contrastive learning enhance foveal attention? (RQ2)
@@
5.3 Does return-guided contrastive learning help the agent focus on task-relevant regions in the presence of distractor objects? (RQ3)
To further investigate whether contrastive learning helps the attention mechanism to filter irrelevant visual information, we train and evaluate on cluttered variants of a subset of the tasks. As shown in the bottom row of Figure 2, the performance gap is more apparent. For example, while foveal attention without contrastive learning is unable to solve the PushTClutter task, even underperforming the baseline, contrastive learning provides the necessary guidance to find critical cues from the cluttered environment. Also for PokeCubeClutter, the foveal attention with contrastive learning attains 50% success rate with the fewest environment steps.
@@ 5.4 Is our approach applicable across different RL algorithms? (RQ4)
Figure 4: SAC results on ManiSkill3 environments: foveal attention with return-guided contrastive learning, CURL, and the SAC baseline. Across most manipulation tasks, return-guided foveal attention provides the most consistent performance improvements.
@@
5.5 How does our approach compare against existing contrastive representation learning baselines? (RQ5)
We compare our method to CURL laskin2020curl, a widely used contrastive representation learning approach that has shown strong performance on visually simple domains such as the DMControl Suite tassa2018deepmind. CURL’s random-crop augmentation encourages global translation invariance and regularizes the encoder effectively. However, prior work has noted that augmentation-based methods can face challenges such as training instability hansen2021stabilizing, sensitivity to augmentation types yarats2021mastering, and difficulty in focusing on task-relevant features bertoin2022look. We hypothesize that these issues may also occur in manipulation settings, where task-relevant cues are highly localized and depend on fine-grained geometry such as object–gripper alignment or contact geometry.
In our experiments (Figure 4), we observe that CURL performs competitively on certain tasks (e.g., PlaceSphere, PokeCube), while less effectively on others. We hypothesize that this variation relate to differences in camera configuration, object scale, or scene complexity of the task rather than any inherent limitation of CURL itself. Across the same tasks, our return-guided foveal attention shows more consistent improvements over the baseline, suggesting that its inductive bias can align well with certain manipulation settings. We emphasize that these observations are empirical and highlight only that different inductive biases may interact differently with the structure and visual requirements of each task. Our implementation of CURL is detailed in Appendix C.
Although both CURL and our approach both use contrastive objectives, they emphasize fundamentally different inductive biases. While CURL promotes global invariance through augmentation-based alignment, our “Gaze on the Prize” objective ties feature similarity to task signals and attentional relevance. These approaches are complementary and the findings suggest exploring hybrid strategies that combine augmentation-based representation learning with reward-guided attention objectives is an interesting direction for future work.
@@ 5.6 Ablations
We conduct ablations on the PushT task under PPO, as this setting shows the largest gap between the foveal attention and our full method with contrastive learning. This makes it the most informative testbed for isolating the effects of each design choice of our method. We focus our ablation on two core hyperparameters that characterize the general contrastive learning pipeline: (1) buffer size, which determines the diversity and temporal coverage of samples available for mining positive and negative pairs, and (2) update frequency, which affects both the computational cost and the consistency of corrective attention signals.
Figure 5: Ablation study on different contrastive buffer sizes.
@@
Buffer size: Contrastive buffer size is an important design choice of our method, as PPO typically uses only the most recent rollout data unlike off-policy algorithms. Prior work has shown that augmenting on-policy algorithms with replay buffers can improve performance wang2016sample; novati2019remember, motivating our use of a separate persistent buffer for contrastive learning. As shown in Fig. 5, using only the on-policy batch ($\approx$20k) still provides improvement over the no-contrastive baseline, demonstrating that even short-horizon features can meaningfully guide the foveal attention. Increasing the buffer to 100k yields the best performance, suggesting that a moderate amount of temporal diversity is beneficial for mining informative triplets. Interestingly, a much larger buffer (200k) still outperforms the no-contrastive variant but shows slower early learning, likely due to mining from stale representations that no longer match the current encoder. This highlights a key trade-off where the buffer must be large enough to provide diverse samples, but not to a degree where outdated features can weaken the quality of contrastive supervision.
Figure 6: Ablation study on contrastive update frequency.
Contrastive Frequency Because contrastive objective adds computational overhead, we evaluate different contrastive update frequencies to check whether reducing the update frequency preserves performance benefits. As shown in Figure 6, performance improves consistently as contrastive updates become more frequent. Updating contrastive learning every iteration achieves the fastest and most stable convergence, while updating every 2 iterations still yields better performance compared to the no-contrastive baseline. In contrast, updating every 4 iterations slows down learning and increases variance. These results indicate that attention benefits from regular corrective signals and if the contrastive objective is applied less frequently, the learned attention can drift and lose task relevance.
@@
Table 1: Throughput reduction and performance trade-offs. Steps Per Second (SPS) is measured during training on PushT task using a single NVIDIA A6000 GPU, representing wall-clock throughput. Sample efficiency is measured as the ratio of steps required to reach 50% success rate (SR) compared to baseline. Detailed breakdown it provided in Appendix B.